OpenStudy (anonymous):

Help, please? :) Which equation is not equivalent to the formula d = rt? A r=t/d B r=d/t C t=d/r D tr

OpenStudy (anonymous):

I believe that the last one is d=tr .. ? Suppose you have a=bc, then a=bc => a=cb => c=a/b => b=a/c And, many more such equations!

OpenStudy (anonymous):

for d=rt the following equations work: r=d/t t=d/r d=tr So that means A. r=t/d is not equivalent. Why? Because to find r you must divide t from both sides making d/t not t/d. Hope this helps! (: